Magnet A has thrice the magnetic field strength of magnet B and at a certain distance pulls on magnet B with a force of 100 Newtons. With how much force, then, does magnet B pull on magnet A? What is the force if the distance between the magnets is doubled?
1
Expert's answer
2020-07-06T15:12:30-0400

(a) Magnets A and B are pulling equally on each other (Newton's Third Law). if A pulls on B with 100 N, then B also pulls on A with 100 N.


(b) The force of interaction between the magnets is inversely proportional to the fourth power of the distance between them, which is a consequence of the dipole-dipole nature of the interaction (for non-existent point magnetic charges, this force would, as usual, be inversely proportional to the square of the distance).

So,


"100=\frac{k}{r^4}" and "f=\frac{k}{(2r)^4}" "\to" "100\cdot r^4=f\cdot 2^4\cdot r^4 \to f=\frac{100}{2^4}=6.25 N"